a circle has a diameter of 4'3 9/16" what is the area of the circle in square inches? what is the area in square feet?
a=pie(don't know where sign is on computer) a = pie x r2
r2= 51.5625 r = 25.78125in2 pie x 25.78125 in2 = 2087.07in2
so area of circle in2 = 2807.07in2?
area of circle in square feet = 28.9870ft2?
hope you can understand that! Thanks alot Tammy

Answer by stanbon(48545)

a circle has a diameter of 4'3 9/16" what is the area of the circle in square inches? what is the area in square feet?
Area = pi x r^2
diameter = 51.5625 in
radius = 25.78125 in ; r^2 = 664.67
so area of circle = (pi)r^2 = 2088.12 sq. in.
area of circle in square feet = (2088.12 sq. in.)/(12^2 sq. in) = 14.5 sq ft
Cheers,
Stan H.